Find a binary function * on Q such that phi is an isomorphism mapping <Q, +> with <Q , *) and give the identity element for *. The map is phi : Q --> Q is given by phi(x)=3x-1 for x is an element of Q


1
Expert's answer
2021-02-24T07:29:27-0500

Solution:

Given an isomorphism mapping, i.e., one-to-one and onto map "\phi: {Q} \rightarrow {Q}" defined as "\phi(x)=3 x-1" .

Note that for any "x \in {Q}" , we have "\phi\left(\frac{x+1}{3}\right)=x"

Here we find out what "a * b" equals.

From the above line, and from the fact that the condition "\phi(x) * \phi(y)=\phi(x+y)" must hold, we have,

"a * b=\phi\left(\frac{a+1}{3}\right) * \phi\left(\frac{b+1}{3}\right)=\phi\left(\frac{a+1}{3}+\frac{b+1}{3}\right)=\phi\left(\frac{a+b+2}{3}\right)=3( \frac{a+b+2}{3})-1=a+b+1"

Therefore our binary operation is "a * b=a+b+1" .

For this particular binary operation the element "-1 \in {Q}" is the identity because "-1 * a=-1+a+1=a" .